Calculate the pH of an aqueous ammonia solution that has an OH-concentration of 3.9 × 103M. Is the solution more basic or acidic


1
Expert's answer
2021-05-18T07:48:54-0400

Solution.

"\begin{bmatrix}
 OH^- \\

\end{bmatrix}=3.9\sdot 10^3M;"

"\begin{bmatrix}
H^+ \\

\end{bmatrix}\begin{bmatrix}
OH^- \\

\end{bmatrix}=1\sdot10^{-14}\implies \begin{bmatrix}
H^+ \\

\end{bmatrix}=\dfrac{1\sdot10^{-14}}{\begin{bmatrix}
 OH^- \\
 
\end{bmatrix}};"

"\begin{bmatrix}
H^+ \\

\end{bmatrix}=\dfrac{1\sdot10^{-14}}{
 3.9\sdot10^3 }=2.56\sdot10^{-18};"

"pH=-log\begin{bmatrix}
H^+ \\

\end{bmatrix}" ;

"pH=-log(2.56\sdot10^{-18})=17.59;"

The solution is basic.

Answer: "pH=17.59,"  the solution is basic.